OpenStudy (anonymous):

evaluate the difference quotient for the given function f(x)= 1/x, f(x)-f(a)/x-a

OpenStudy (anonymous):

\[f(x)= 1/x\]\[f(a) = 1/2\] \[\frac{f(x)-f(a)}{x-a} =\frac{\frac 1x -\frac 1a}{x-a} \]

OpenStudy (anonymous):

You have to simplify it more.

OpenStudy (anonymous):

What do I do next?

OpenStudy (anonymous):

First simplify the numerator: \[ \frac 1 x - \frac 1 a = ? \]

OpenStudy (anonymous):

\[ \frac 1x - \frac 1a = \frac a{xa}-\frac x{xa} = \frac{a-x}{xa} \]

OpenStudy (anonymous):

So we have \[\large \frac{\frac{a-x}{xa}}{x-a} = \frac{a-x}{(xa)(x-a)} \]

OpenStudy (anonymous):

and so the x-a and a-x cancels out?

OpenStudy (anonymous):

Yeah, they cancel out to \(-1\).

OpenStudy (anonymous):

Assuming that \(x\neq a\)

OpenStudy (anonymous):

thank you!